OpenStudy (anonymous):

How can I solve the following system of equations? g + 25 = s g + s = 425

OpenStudy (dtan5457):

I would isolate g on the 2nd equation So you have g+25=s and g=425-s Plug that back into the first equation

OpenStudy (anonymous):

So far, this is what I got: g + (g + 25) = 425

OpenStudy (anonymous):

Oh! Okay! :)

OpenStudy (dtan5457):

If you use my technique you should have 425-s+25=s solve for s. Get that value to solve for g by plugging it back into either equation

OpenStudy (anonymous):

Yep, I think I have got it from here.

OpenStudy (dtan5457):

Awesome. Feel free to let me know what you get so I can check for you.

OpenStudy (anonymous):

Okay! Thanks!

OpenStudy (anonymous):

Hm. Maybe not. What would the next step be?

OpenStudy (anonymous):

To isolate s on the left side?

OpenStudy (anonymous):

@dtan5457 ^

OpenStudy (anonymous):

wait do you still need help?

OpenStudy (anonymous):

Yes.

Nnesha (nnesha):

425-s+25=s combine like terms

OpenStudy (anonymous):

Would we get the s's on the same side?

OpenStudy (anonymous):

hmmmmmm.. sorry........... i dont kno w............................. i dont t hink i learned this stuff yet .......... sorry!!!!!!!!!!!!

Nnesha (nnesha):

yes but you can combine 425 and 25 first

OpenStudy (anonymous):

It's okay! @Saisuke<3

OpenStudy (anonymous):

Okay! Then I would get ...

Nnesha (nnesha):

then add s both side

OpenStudy (anonymous):

400 - s = s ?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Wouldn't we subtract s from both sides?

Nnesha (nnesha):

425-s+25=s 425 + 25 =??

OpenStudy (anonymous):

425 + 25 = 450

Nnesha (nnesha):

yes so now you have add s both side